OpenStudy (anonymous):

Solve by completing the square x^2+16x=-2

OpenStudy (anonymous):

What step do you not understand?

OpenStudy (anonymous):

What value do I need to add to both sides?

OpenStudy (anonymous):

(B/2)^2

OpenStudy (anonymous):

Look at it like this: you're creating a square which GIVES you more than you need, so you need to remove that scruff.

OpenStudy (anonymous):

So if you have x^2+b+c=0, the square you want will be (x+b/2)^2, but (x+b/2)^2=x^2+bx+(b/2)^2 (x+b/2)^2=x^2+bx+ --->(b/2)^2<--- THIS, you need to remove in order for you to be able to make the square.

OpenStudy (anonymous):

16/2=8 8^2=64. You'll add 64 to each side

OpenStudy (anonymous):

x^2+16x+64=-2+64 (X+8)^2=62 X+8=pos or neg sqrt62 X= -8 + sqrt62 or x =-8-sqrt62